### v2

@@ 4 Summary and conclusions
We have developed a novel robust numerical method that allows us to obtain the steady-state force-free dissipationless ideal 3D pulsar magnetosphere on a standard off-the-shelf computer with a GPU. Our method does not involve a time-dependent simulation that eventually relaxes to a steady-state, as in all previous attempts to obtain the steady-state pulsar magnetosphere. Instead, we solve directly the generalization of the axisymmetric pulsar equation (Scharlemann & Wagoner, 1973) in 3D (Endean, 1974; Mestel, 1975). Notice that the method is not limited to centrally located dipole fields and can accomodate any stellar magnetic field configuration.
@@
Our methodology involves the decomposition of the pulsar magnetosphere into the regions of open and closed field lines. This allows us to treat the separatrix current sheet as a perfect mathematical discontinuity between the two regions. We also ‘numerically remove’ the equatorial current sheet by reversing the field direction in the southern hemisphere, and reverting it to its true direction after the 3D solution is obtained. Our present method can be directly applied to more general ideal MHD problems that contain current sheets, as for example in active regions of the solar corona. In this paper we present our first solution for one particular size and shape (circular) of the polar cap of open magnetic field lines, and for one particular pulsar inclination angle with respect to the axis of rotation. The solution is obtained in a coordinate system that corotates with the star. Our results elucidate the potential of our method to generate the reference solution against which we will compare with current numerical solutions of the pulsar magnetosphere.
@@
In this paper, we have trained a set of two NNs that instantaneously yields the three components $B_{x},B_{y},B_{z}$ of the magnetic field and their spatial derivatives at any given point $(x,y,z)$ of the magnetosphere for one particular pulsar inclination, polar cap size and shape. This is extremely helpful if one wants to draw magnetospheric features like magnetic and electric field lines, electric charge and electric current distributions, Poynting flux, current sheets, etc. We will present a detailed investigation of the pulsar magnetosphere for various pulsar inclinations and for various polar cap sizes and shapes in an extended forthcoming publication, and we will produce corresponding trained NNs (Paper IV). We are anxious to learn whether magnetospheric solutions exist even for arbitrarily small polar caps with arbitrarily small luminosities as seen in 2D (Paper II). We would like to emphasize that our methodology allows us to obtain new solutions of the ideal 3D pulsar magnetosphere by determining a priori the shape and size of the polar cap, something that current conventional methods are unable to do. Such solutions were never obtained before because all numerical solutions obtained in the literature always relax to one particular magnetospheric configuration, i.e. to one particular polar cap shape and size determined by the simulation, without any possibility for control by the programmer. This opens up a whole spectrum of solutions between which the real pulsar magnetosphere may choose to transition during its evolution. More work is needed along these lines.

### v3

@@ 4 Summary and conclusions
We have developed a novel robust numerical method that allows us to calculate steady-state force-free ideal 3D pulsar magnetospheres for various pulsar inclinations and polar cap shapes. Our method does not involve a time-dependent simulation that eventually relaxes to a steady-state, as in all previous attempts to obtain the steady-state pulsar magnetosphere. Instead, we solve directly the generalization of the axisymmetric pulsar equation (Scharlemann & Wagoner, 1973) in 3D (Endean, 1974; Mestel, 1975). Notice that the method is not limited to centrally located dipole fields and can accomodate any stellar magnetic field configuration. For example, one may determine a particular magnetic field configuration with hotspots that reproduce the NICER X-ray observations, assume that these hotspots correspond to the origin of open magnetic field lines (i.e. assume that these are the polar caps), and calculate the 3D ideal magnetosphere that corresponds to these particular polar caps using our methodology.
@@
Our methodology involves the decomposition of the pulsar magnetosphere into the regions of open and closed field lines. This allows us to treat the separatrix current sheet as a perfect mathematical discontinuity between the two regions. We also ‘numerically remove’ the equatorial current sheet by reversing the field direction in the southern hemisphere, and reverting it to its true direction after the 3D solution is obtained. We do acknowledge that although this approach helps in dealing with magnetospheric current sheets numerically, it completely ignores the physics of reconnection, electromagnetic energy dissipation, plasma instabilities and particle acceleration along them. In that sense, our solutions may be considered only as ideal solutions that must be augmented with the microphysics of the current sheets. Our present method can be directly applied to more general ideal MHD problems that contain current sheets, as for example in active regions of the solar corona. In this paper we present our first solution for one particular size and shape (circular) of the polar cap of open magnetic field lines, and for one particular pulsar inclination angle with respect to the axis of rotation. The solution is obtained in a coordinate system that corotates with the star. Our results elucidate the potential of our method to generate new solutions of the ideal pulsar magnetosphere.
@@
In this paper, we have trained a set of two NNs that instantaneously yields the three components $B_{x},B_{y},B_{z}$ of the magnetic field and their spatial derivatives at any given point $(x,y,z)$ of the magnetosphere for one particular pulsar inclination, and one particular polar cap size and shape. This is extremely helpful if one wants to draw magnetospheric features like magnetic and electric field lines, electric charge and electric current distributions, Poynting flux, current sheets, etc. We will present a detailed investigation of the pulsar magnetosphere for various pulsar inclinations and for various polar cap sizes and shapes in an extended forthcoming publication, and we will produce corresponding trained NNs (Paper IV). We look forward to learn whether magnetospheric solutions exist even for arbitrarily small polar caps with arbitrarily small luminosities as seen in 2D (Paper II). We would like to emphasize that our methodology allows us to obtain new solutions of the ideal 3D pulsar magnetosphere by determining a priori various shapes and sizes of the polar cap. Such solutions were never obtained before because all numerical solutions obtained in the literature always relax to one particular magnetospheric configuration, i.e. to one particular polar cap shape and size determined by the simulation, without any possibility for control by the programmer. This opens up a whole spectrum of solutions between which the real pulsar magnetosphere may choose to transition during its evolution. More work is needed along these lines.